# Geometry of exactness of moment-SOS relaxations for polynomial optimization

### 01 Dec 2023 v2

@@ 1 Solving POPs with the moment-SOS hierarchy
$\begin{array}[]{rcllcll}v^{r}(f)&:=&\inf_{{\bf y}}&\ell_{{\bf y}}(f)&\geq&\sup_{v}&v\\ &&\mathrm{s.t.}&{\bf y}\in{\mathscr{R}}({\bf g})^{r}_{d}&&\mathrm{s.t.}&f-v\in{\mathscr{Q}}({\bf g})^{r}_{d}\\ &&&\ell_{{\bf y}}(1)=1.\end{array}$ (3)
This primal-dual pair of semidefinite optimization problems is called the moment-SOS relaxation of order $r$
@@
Since ${\mathscr{X}}$ is bounded, it is always possible to add a redundant quadratic constraint $R^{2}-\sum_{k=1}^{n}x^{2}_{i}\geq 0$ to the description of ${\mathscr{X}}$. So Assumption 1 is without loss of generality. Then it follows from [6] that the relaxed values are monotonically converging lower bounds on the value:
$v^{r}(f)\leq v^{r+1}(f)\leq v^{\infty}(f)=v(f),$ (4)
that the primal is attained (i.e. the infimum is a minimum) and that there is no duality gap (i.e. the infimum equals the supremum). Moreover if ${\mathscr{X}}$ has an interior point, then the dual is attained (i.e. the supremum is a maximum). If ${\mathscr{X}}$ does not have an interior point, e.g. if it is a low-dimensional algebraic variety, then additional algebraic or geometric conditions are required for the dual to be attained, cf. [2, 1, 8].
@@ 2 Exactness cone
Beyond the asymptotic convergence guarantees (4), it is important to know whether the moment-SOS relaxation of a given order is exact, i.e. whether an optimal solution to the relaxation attains the value $v(f)$. If this is the case, there is no need to increase $r$ and solve larger semidefinite optimization problems.
@@ 3 Main result
Proof: At order $r$, exactness of the moment-SOS relaxation (3) is equivalent to the existence of an optimal primal-dual pair ($\hat{{\bf y}},\hat{v}$) with a Dirac vector $\hat{{\bf y}}\in{\mathscr{R}}({\bf g})^{r}_{d}$ concentrated at a point $\hat{{\bf x}}\in{\mathscr{X}}$, i.e. such that the objective function is the point evaluation $v^{r}(f)=\ell_{\hat{{\bf y}}}(f)=f(\hat{{\bf x}})=\hat{v}$, and also $f-\hat{v}\in{\mathscr{Q}}({\bf g})^{r}_{d}$. Since $\hat{{\bf x}}\in{\mathscr{X}}$, the value $f(\hat{{\bf x}})$ is an upper bound on the global optimum $v(f)$. According to (4), $v^{r}(f)$ is also a lower bound, so it follows that $v^{r}(f)=v(f)$.
Optimality of a primal-dual pair $(\hat{{\bf y}},\hat{v})$ is equivalent to conic complementarity of $\hat{{\bf y}}\in{\mathscr{R}}({\bf g})^{r}_{d}$ and $f-\hat{v}\in{\mathscr{Q}}({\bf g})^{r}_{d}$, i.e. $\ell_{\hat{{\bf y}}}(f)-\hat{v}=f(\hat{{\bf x}})-\hat{v}=f(\hat{{\bf x}})-v(f)=0$. Therefore, for each $\hat{{\bf x}}\in{\mathscr{X}}$, objective functions $f\in\mathbb{R}[{\bf x}]_{d}$ that achieve exactness are such that $f=v(f)+f^{0}$ for some polynomial $f^{0}\in{\mathscr{Q}}({\bf g})^{r}_{d}$. This polynomial satisfies $f^{0}(\hat{{\bf x}})=f(\hat{{\bf x}})-v(f)=0$, i.e. it belongs to ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$.
Cone ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ is semidefinite representable since it is the projection of linear sections of the SOS cone $\Sigma[{\bf x}]$, which is itself semidefinite representable.
The proof is completed by observing that ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ reduces to the origin whenever $\hat{{\bf x}}$ is not optimal for POP (1). Let us proceed by contradiction. Let $\hat{{\bf x}}\in{\mathscr{X}}$ be non-optimal, and assume there exists a non-zero polynomial $f^{0}\in{\mathscr{S}}_{\hat{{\bf x}}`}({\bf g})^{r}_{d}$, i.e. $f^{0}(\hat{{\bf x}})=0$. Since $\hat{{\bf x}}$ is not optimal, $f^{0}(\hat{{\bf x}})=f(\hat{{\bf x}})-v(f)>0$, a contradiction. $\Box$
@@
In our definition of the exactness cone, we require that both primal and dual values are attained in (3). Since what matters is whether $v^{r}(f)=v(f)$, we may relax our attainment requirements. The corresponding exactness cones would be slightly larger, at the price of more technicalities.
@@ 5.1 Finite set
$\begin{array}[]{rcl}{\mathscr{S}}(2,2)^{1}_{1}&=&\{f\in\mathbb{R}[{\bf x}]_{1}:f({\bf x})=f_{0}+f_{1}x_{1}+f_{2}x_{2}=s_{0}({\bf x})+s_{1}g_{1}({\bf x})+s_{2}g_{2}({\bf x}),\\ &&\quad f(2,2)=0,s_{0}\in\Sigma[{\bf x}]\cap\mathbb{R}[{\bf x}]_{2},s_{1}\in\mathbb{R},s_{2}\in\mathbb{R}\}\end{array}$
which can be written more explicitly by expressing the quadratic SOS constraint
@@
with a 3-by-3 positive semidefinite Gram matrix $X$, as well as the equality constraint
$f(2,2)=s_{0}(2,2)=x_{11}+4x_{21}+4x_{22}+4x_{31}+8x_{32}+4x_{33}=0.$
@@
$\begin{array}[]{rcl}{\mathscr{S}}(2,2)^{1}_{1}&=&\{(f_{0},f_{1},f_{2})\in\mathbb{R}^{3}:f_{0}=x_{11},f_{1}=2x_{21}-s_{2},f_{2}=2x_{31}+2s_{1}+s_{2},\\ &&\quad 0=x_{22}+s_{2},0=2x_{32}+s_{1},0=x_{33}-2s_{1}-s_{2},\\ &&\quad x_{11}+4x_{21}+4x_{22}+4x_{31}+8x_{32}+4x_{33}=0,\\ &&\quad X\succeq 0,s_{1}\in\mathbb{R},s_{2}\in\mathbb{R}\}\\ &&=\{(f_{0},f_{1},f_{2})\in\mathbb{R}^{3}:\left(\begin{array}[]{ccc}2f_{0}&f_{1}+s_{2}&f_{2}-2s_{1}+s_{2}\\ f_{1}+s_{2}&-2s_{2}&-s_{1}\\ f_{2}-2s_{1}+s_{2}&-s_{1}&4s_{1}+2s_{2}\end{array}\right)\succeq 0,s_{1}\in\mathbb{R},s_{2}\in\mathbb{R}\}\end{array}$
is the 3-dimensional projection of a 5-dimensional cubic spectrahedral cone.
@@
On Figure 2 we represent the same objects for the second relaxation, i.e. $r=2$. On the left, we see that the moment relaxation ${\mathscr{R}}({\bf g})^{2}_{1}$ is the polytope $\mathrm{conv}\>{\mathscr{X}}$, i.e. ${\mathscr{R}}({\bf g})^{2}_{1}\setminus\mathrm{conv}\>{\mathscr{X}}$ is empty: the tiny dark gray regions of Figure 1(a) disappeared. Consistently with Lemma LABEL:polytope, we observe on the right that the exactness cone ${\mathscr{F}}({\bf g})^{2}_{1}$ is the whole space $\mathbb{R}[{\bf x}]_{1}$, i.e. the relaxation is exact everywhere: the tiny white regions of Figure 1(b) disappeared, consistently with Lemma 1. Exactness follows from the property that all non-negative bivariate quartics are SOS. Indeed, the dual problem consists of maximizing $v$ such that $f-v$ is positive on the four points of ${\mathscr{X}}$, a linear constraint. But this is equivalent to enforcing that $f-v$ is a degree 4 (i.e. $r=2$) SOS subject to the linear constraint.

### 28 Feb 2024 v3

@@ 1 Solving POPs with the moment-SOS hierarchy
$\begin{array}[]{rcllcrcll}\text{mom}(f)^{r}&:=&\inf_{{\bf y}}&\ell_{{\bf y}}(f)&\geq&\text{sos}(f)^{r}&:=&\sup_{v}&v\\ &&\mathrm{s.t.}&{\bf y}\in{\mathscr{R}}({\bf g})^{r}_{d}&&&&\mathrm{s.t.}&f-v\in{\mathscr{Q}}({\bf g})^{r}_{d}\\ &&&\ell_{{\bf y}}(1)=1.\end{array}$ (3)
This primal-dual pair of semidefinite optimization problems is called the moment-SOS relaxation of order $r$. Note that by construction
$\text{sos}(f)^{r}\leq\text{mom}(f)^{r}\leq v(f).$ (4)
@@
Since ${\mathscr{X}}$ is bounded, it is always possible to add a redundant quadratic constraint $R^{2}-\sum_{k=1}^{n}x^{2}_{i}\geq 0$ to the description of ${\mathscr{X}}$. So Assumption 1 is without loss of generality. Then it follows from [6] that in (3) the primal is attained (i.e. the infimum is a minimum), there is no duality gap (i.e. the infimum equals the supremum) and the relaxed values are monotonically converging lower bounds on the value:
$\text{sos}(f)^{r}=\text{mom}(f)^{r}\leq\text{sos}(f)^{r+1}=\text{mom}(f)^{r+1}\leq\text{sos}(f)^{\infty}=\text{mom}(f)^{\infty}=v(f).$
Moreover if ${\mathscr{X}}$ has an interior point, then the dual is attained (i.e. the supremum is a maximum). If ${\mathscr{X}}$ does not have an interior point, e.g. if it is a low-dimensional algebraic variety, then additional algebraic or geometric conditions are required for the dual to be attained, cf. [2, 1, 8].
@@ 2 Exactness cone
Beyond asymptotic convergence guarantees, it is important to know whether the moment-SOS relaxation of a given order $r$ is exact, i.e. whether $\text{sos}(f)^{r}=v(f)$. If this is the case, there is no need to increase $r$ and solve larger semidefinite optimization problems.
@@ 3 Main result
Proof: If $f\in{\mathscr{F}}({\bf g})^{r}_{d}$ then $\text{sos}(f)^{r}=v(f)=f(\hat{{\bf x}})$ where $\hat{{\bf x}}$ is a global minimizer of $f$ on ${\mathscr{X}}$. As $\text{sos}(f)^{r}$ is attained, we have $f-\text{sos}(f)^{r}=f-f(\hat{{\bf x}})\in{\mathscr{Q}}({\bf g})^{r}_{d}$ and thus $f$ belongs to ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ for some $\hat{{\bf x}}\in{\mathscr{X}}$.
Conversely, assume $f\in{\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ for some $\hat{{\bf x}}\in{\mathscr{X}}$. Then from (4) we have $f(\hat{{\bf x}})\leq\text{sos}(f)^{r}\leq\text{mom}(f)^{r}\leq v(f)\leq f(\hat{{\bf x}})$, which implies that $\text{sos}(f)^{r}=\text{mom}(f)^{r}=v(f)$ and hence $f\in{\mathscr{F}}({\bf g})^{r}_{d}$.
Cone ${\mathscr{S}}_{\hat{{\bf x}}}({\bf g})^{r}_{d}$ is semidefinite representable since it is the projection of linear sections of the SOS cone $\Sigma[{\bf x}]$, which is itself semidefinite representable. $\Box$
@@
[…]
@@ 5.1 Finite set
$\begin{array}[]{rcl}{\mathscr{S}}(2,2)^{1}_{1}&=&\{f\in\mathbb{R}[{\bf x}]_{1}:f({\bf x})-f(2,2)=s_{0}({\bf x})+s_{1}g_{1}({\bf x})+s_{2}g_{2}({\bf x}),\\ &&\quad s_{0}\in\Sigma[{\bf x}]\cap\mathbb{R}[{\bf x}]_{2},s_{1}\in\mathbb{R},s_{2}\in\mathbb{R}\}\end{array}$
where $f({\bf x})=f_{0}+f_{1}x_{1}+f_{2}x_{2}$, which can be written more explicitly by expressing the quadratic SOS constraint
@@
with a 3-by-3 positive semidefinite Gram matrix $X$, and identifying like powers of ${\bf x}$ in the equation $f({\bf x})-f(2,2)=s_{0}({\bf x})+s_{1}g_{1}({\bf x})+s_{2}g_{2}({\bf x})$:
$\begin{array}[]{rcl}-2f_{1}-2f_{2}&=&x_{11}\\ f_{1}&=&2x_{21}-s_{2}\\ f_{2}&=&2x_{31}+2s_{1}+s_{2}\\ 0&=&x_{22}+s_{2}\\ 0&=&2x_{32}+s_{1}\\ 0&=&x_{33}-2s_{1}-s_{2}.\end{array}$
@@
${\mathscr{S}}(2,2)^{1}_{1}=\mathbb{R}\oplus\{(f_{1},f_{2})\in\mathbb{R}^{2}:\left(\begin{array}[]{ccc}-4f_{1}-4f_{2}&\star&\star\\ f_{1}+s_{2}&-2s_{2}&\star\\ f_{2}-2s_{1}-s_{2}&s_{1}&4s_{1}+2s_{2}\end{array}\right)\succeq 0,(s_{1},s_{2})\in\mathbb{R}^{2}\}$
is a projection of a 4-dimensional cubic spectrahedral cone.
@@
On Figure 2 we represent the same objects for the second relaxation, i.e. $r=2$. On the left, we see that the moment relaxation ${\mathscr{R}}({\bf g})^{2}_{1}$ is the polytope $\mathrm{conv}\>{\mathscr{X}}$, i.e. ${\mathscr{R}}({\bf g})^{2}_{1}\setminus\mathrm{conv}\>{\mathscr{X}}$ is empty: the tiny dark gray regions of Figure 1(a) disappeared. We observe on the right that the exactness cone ${\mathscr{F}}({\bf g})^{2}_{1}$ is the whole space $\mathbb{R}[{\bf x}]_{1}$, i.e. the relaxation is exact everywhere: the tiny white regions of Figure 1(b) disappeared, consistently with Lemma 1. Exactness follows from the property that all non-negative bivariate quartics are SOS. Indeed, the dual problem consists of maximizing $v$ such that $f-v$ is positive on the four points of ${\mathscr{X}}$, a linear constraint. But this is equivalent to enforcing that $f-v$ is a degree 4 (i.e. $r=2$) SOS subject to the linear constraint.
@@ 5.2 Non-convex set
Acknowledgement
The statement of Theorem 1 and its proof were given by an anonymous reviewer.